# Part I: Nonlinear Codes

### v1

@@ III Direct Constructions for de Bruijn Arrays Codes
Let ${\cal F}$ be a PF$(n,k)$, $m=2^{t}-1$, $t\geq k$, and define the following set of matrices
$\mathbb{C}\triangleq\{[X_{i_{1}},{\bf E}^{j_{2}}X_{i_{2}},\ldots,{\bf E}^{j_{m}}X_{i_{m}},{\bf E}^{j_{m+1}}X_{i_{m+1}}]~{}:~{}$
$0\leq j_{r}\leq 2^{k}-1,~{}2\leq r\leq m+1\},$
where each $X_{i_{r}}$ is a column vector and its cyclic shifts is related to its zero state. Then, the code $\mathbb{C}$ is a $(2^{k},2^{t};n,m)$-DBAC of size $2^{nm-k-t}$.
@@
The codewords of the construction of $\mathbb{C}$ are $2^{k}\times 2^{t}$ matrices, but they can be also viewed as cycles of length $2^{t}$ in the graph $G_{{\cal F},m}$. We start by showing that the horizontal period of each cycle defined in $\mathbb{C}$ is ${m+1=2^{t}}$, i.e., the associated matrix does not have a smaller horizontal periodicity. If the period of some cycle ${\cal C}$ in $\mathbb{C}$ is smaller than $2^{t}$, i.e., the associated matrix has a smaller horizontal periodicity, then it should be a divisor of $2^{t}$, i.e., a power of 2. Assume for the contrary that such a matrix has a horizontal periodicity, i.e., it has period $2^{s}$ for some $s$, ${0\leq s<t}$ which implies that it consists of $2^{s}$ disjoint $2^{k}\times 2^{t-s}$ equal vertically periodic matrices. Since $\sum_{r=1}^{2^{t}}i_{r}\equiv 1~{}({\mbox{mod}}~{}2^{n-k})$, it follows that $\sum_{r=1}^{2^{t}}i_{r}=\alpha 2^{n-k}+1$, for some positive integer $\alpha$, and hence $\sum_{r=1}^{2^{s}}i_{r}=\frac{\alpha 2^{n-k}+1}{2^{t-s}}$. But, $\frac{\alpha 2^{n-k}+1}{2^{t-s}}$ is not an integer, a contradiction. Therefore, all the cycles of $\mathbb{C}$ have no smaller horizontal periodicity and hence they form $2^{k}\times 2^{t}$ arrays with period $2^{t}$.
We continue by computing the size of the code $\mathbb{C}$. Each $X_{i_{r}}$, $1\leq r\leq m$, is taken arbitrarily from ${\cal F}$ and hence it can be chosen in $2^{n-k}$ distinct ways. The sequence $X_{i_{m+1}}$ is determined by the equation ${\sum_{r=1}^{m+1}i_{r}\equiv 1~{}({\mbox{mod}}~{}2^{n-k})}$. The first cycle $X_{i_{1}}$ is taken in its zero state, while the next $m-1$ cycles can be taken in any of their $2^{k}$ possible cyclic shifts. The shift of the last cycle $X_{i_{m+1}}$ is determined by the equation $\sum_{r=2}^{m+1}j_{r}\equiv 0~{}({\mbox{mod}}~{}2^{k})$. Hence, each one of the middle $m-1$ cycles can be chosen in $2^{n-k}2^{k}=2^{n}$ distinct ways. The first cycle can be chosen in $2^{n-k}$ distinct ways and the last cycle and its shift are determined by the first $m$ cycles and their shifts. Therefore, there are together $2^{n(m-1)}2^{n-k}=2^{nm-k}$ distinct ways to choose all the cycles and their shifts.
Now, we calculate the number of cycles which are counted more than once in this enumeration. Each constructed cycle ${\cal C}$ can start with any one of its $m+1=2^{t}$ columns. Such a column, say the $r$-th column, will be taken in the shift where its zero state is the first $n$-tuple. The other columns are defined by the order of the columns in ${\cal C}$ and their shifts are taken exactly as in ${\cal C}$ related to the first state of the new first column. Therefore, each codeword ${\cal C}\in\mathbb{C}$ is constructed exactly $2^{t}$ times in the construction of $\mathbb{C}$. Thus, the size of $\mathbb{C}$ is $2^{nm-k}/2^{t}=2^{nm-k-t}$. Note, that the sum of all the shifts, in this new arrangement of the first column, will remain congruent to 0 modulo $2^{k}$ since $t\geq k$ and hence the sum of the new shifts will be changed by $j_{r}2^{t}$ modulo $2^{k}$ which is congruent to 0 modulo $2^{k}$, where the $r$-th column of ${\cal C}$ is the new first column.
Therefore, the number of $n\times m$ windows in the all the codewords of $\mathbb{C}$ is $2^{nm-k-t}\cdot 2^{k}\cdot 2^{t}=2^{nm}$. Hence, to complete the proof it is sufficient to show that each $n\times m$ matrix appears as a window in a codeword of $\mathbb{C}$. Consider such an $n\times m$ matrix $A$. The columns of $A$ are $n$-tuples in cycles of ${\cal F}$ and hence these cycles can be arranged in a codeword ${\cal C}\in\mathbb{C}$ with all the necessary shifts as required by $A$. This arrangements can be guaranteed by the definition of the first $m=2^{t}-1$ columns in a codeword of $\mathbb{C}$ to form a window with the matrix $A$ in ${\cal C}$, since there are no constraints on the shifts in the construction for the $m-1$ columns which follows the first one.
Thus, $\mathbb{C}$ is a $(2^{k},2^{t};n,m)$-DBAC of size $2^{nm-k-t}$. ∎

### v2

@@ IV Recursive Constructions for de Bruijn Array Codes
The code $\mathbb{C}^{\prime}$ defined in Construction 5 is an $(2^{k},2^{m};n+1,m)$-DBAC of size $2^{m}\Delta$.
@@
From the discussion in Section II-C, by applying ${\bf D}^{-1}$ on a column from $A$, we obtain two complementary sequences of the same period $2^{k}$. Since $A$ is a $2^{k}\times 2^{m}$ doubly-periodic array, it follows that also $B$ is a $2^{k}\times 2^{m}$ doubly-periodic array.
Since the code $\mathbb{C}$ is of size $\Delta$ and each array $A$ is used to construct $2^{m}$ new codewords by using the $2^{m}$ cyclic shifts of ${\cal S}$, it follows that the new code $\mathbb{C}^{\prime}$ has $2^{m}\Delta$ codewords. Each codeword in $\mathbb{C}^{\prime}$ has $2^{k+m}$ distinct windows of size $(n+1)\times m$, giving a total of $2^{k+m}2^{m}\Delta$ windows of size $(n+1)\times m$ in all arrays of $\mathbb{C}$. Since $2^{k+m}2^{m}\Delta=2^{nm}2^{m}=2^{(n+1)m}$ (note that $2^{k+m}\Delta=2^{nm}$ because $\mathbb{C}$ is a $(2^{k},2^{m};n,m)$-DBAC with $\Delta$ codewords), it follows that to complete the proof it is sufficient to show that all the $(n+1)\times m$ windows in the codewords of $\mathbb{C}^{\prime}$ are distinct.
Assume for the contrary that there exist two distinct $(n+1)\times m$ windows $X$ and $Y$ in codewords of $\mathbb{C}^{\prime}$, $B_{1}$ and $B_{2}$ (not necessarily distinct) obtained from codewords $A_{1}$ and $A_{2}$ of $\mathbb{C}$ (not necessarily distinct), respectively, for which $X=Y$ (if $A_{1}=A_{2}$ then $X$ and $Y$ are windows in distinct positions). Let ${\cal X}$ and ${\cal Y}$ be the two $n\times m$ arrays obtained from $X$ and $Y$, respectively, by applying ${\bf D}$ to the columns of $X$ and $Y$, respectively (note that these columns are acyclic sequences). Clearly, if $X=Y$, then ${\cal X}={\cal Y}$. Since $A_{1}$ and $A_{2}$ are codewords in the $(2^{k},2^{m};n,m)$-DBAC $\mathbb{C}$, it follows that all the $n\times m$ windows in its codewords are distinct and hence ${\cal X}={\cal Y}$ implies that they are obtained from the same position of a codeword $A$, i.e., $A_{1}=A_{2}$, using distinct shifts of ${\cal S}$. Moreover, ${\cal X}={\cal Y}$ implies that they are in columns of $B_{1}$ and $B_{2}$ projected from two sets of $m$ columns associated with the same $m$-tuple in ${\cal S}$. But, this implies that $B_{1}=B_{2}$ and ${\cal S}$ is taken with the same shift to obtain $A$ and hence $X$ and $Y$ are the same $(n+1)\times m$ window with the same position in $A$, a contradiction. Thus, $\mathbb{C}^{\prime}$ is an $(2^{k},2^{m};n+1,m)$-DBAC of size $2^{m}\Delta$. ∎
The next construction is a generalization of Construction 5 and therefore its analysis should be similar.
Construction 6
Let $\mathbb{C}$ be a $(2^{k},2^{m-\ell};n,m)$-DBAC of size $\Delta$, where $\ell\geq 0$, $m<2^{m-\ell}$, whose columns are sequences with even weight and $k,n,m\geq 2$. Let ${\cal F}$ be a PF$(m,m-\ell)$ and $A=[{\cal A}_{1}{\cal A}_{2}~{}\cdots~{}{\cal A}_{2^{m}}]$ be any codeword (a $2^{k}\times 2^{m}$ doubly-periodic array) of $\mathbb{C}$. Let
${\cal S}=[s_{1},s_{2},\ldots,s_{2^{m}}]$
be any sequence of ${\cal F}$.
For the code $\mathbb{C}^{\prime}$ we define a new array $B$. Let the $i$-th column of $B$ be $B_{i}={\bf D}^{-1}_{s_{i}}{\cal A}_{i}$, where $1\leq i\leq 2^{m}$. The same method is applied with the $2^{m-\ell}$ distinct cyclic shifts of ${\cal S}$ to obtain more codewords for $\mathbb{C}^{\prime}$.
Theorem 8
The code $\mathbb{C}^{\prime}$ defined in Construction 6 is an $(2^{k},2^{m-\ell};n+1,m)$-DBAC of size $2^{m}\Delta$.
Proof:
Note first that by Theorem 2 there exists a PF$(m,m-\ell)$. The proof is very similar to the proof of Theorem 7 and hence it is omitted. ∎
Construction 7
Let $\mathbb{C}$ be a $(2^{k},2^{t};n,m)$-DBAC of size $\Delta$, where $k,n,t,m\geq 2$, whose columns are sequences with even weight. Let $A$ be any codeword in $\mathbb{C}$ and define
Let
${\cal S}=[s_{0},s_{1},\ldots,s_{2^{m}-1}]$
be a span $m$ de Bruijn sequence whose first $m$ bits are zeros. Define the sequence
of length $2^{t}+(2^{m}-1)2^{t}=2^{m+t}$ and suppose $B=(b_{1},b_{2},\ldots,b_{2^{m+t}})$. Define the following codeword ${\cal B}$ in a new code $\mathbb{C}^{\prime}$: let the $i$-th column of ${\cal B}$ be ${\cal B}_{i}={\bf D}^{-1}_{b_{i}}{\cal A}_{i}$, where ${\cal A}_{i}$ is the $i$-th column of ${\cal A}$ and $1\leq i\leq 2^{m+t}$.
Theorem 9
The code $\mathbb{C}^{\prime}$ defined in Construction 7 is a $(2^{k},2^{m+t};n+1,m)$-DBAC of size $\Delta$.
Proof:
It is an immediate observation that ${\cal A}$ is an $2^{k}\times 2^{m+t}$ array. If we apply the operator ${\bf D}^{-1}$ to a column from $A$, the outcome is two complementary sequences of the same length $2^{k}$. Note also that the sequence $[s_{1},s_{2},\ldots,s_{2^{m}-1}]$ is the shortened de Bruijn sequence obtained from ${\cal S}$ (starts with $m-1$ zeros).
The codeword ${\cal B}$ is an $2^{k}\times 2^{m+t}$ doubly-periodic array and hence it has $2^{k+m+t}$ windows of size $(n+1)\times m$. Since $2^{k+m}2^{t}\Delta=2^{m}2^{nm}=2^{(n+1)m}$ (since $2^{k}2^{t}\Delta=2^{nm}$), it follows that to complete the proof it is sufficient to show that each $(n+1)\times m$ binary array is contained at most once as an $(n+1)\times m$ window in some ${\cal B}$.
Assume for the contrary that there exist two distinct $(n+1)\times m$ windows ${\cal X}$ and ${\cal Y}$ in two codewords ${\cal B}_{1}$ and ${\cal B}_{2}$ of $\mathbb{C}^{\prime}$ (not necessary distinct) for which ${\cal X}={\cal Y}$. Let $X$ and $Y$ be the $n\times m$ arrays obtained from ${\cal X}$ and ${\cal Y}$, respectively, by applying ${\bf D}$ to the columns of ${\cal X}$ and ${\cal Y}$, respectively (note that these columns are acyclic sequences). Clearly, if ${\cal X}={\cal Y}$, then $X=Y$. Since $A$ is a codeword in an $(2^{k},2^{t};n,m)$-DBAC, it follows that all its $n\times m$ windows are distinct and hence $X=Y$ implies that they are obtained from the same position of $A$ in different locations of ${\cal A}$. Moreover, $X=Y$ implies that ${\cal X}$ and ${\cal Y}$ are in columns of ${\cal B}_{1}$ and ${\cal B}_{2}$ projected from two sets of $m$ columns associated with the same $m$-tuple in $B_{1}$ and $B_{2}$. The fact that g.c.d.$(2^{t},2^{m}-1)=1$ implies that the projection of two sets of $m$ columns on ${\cal A}$ and the associated two sets of $m$ entries in $B_{1}$ and $B_{2}$ (where the two sets of columns from ${\cal A}$ yield the same $2^{k}\times m$ window) cannot yield the same pair of $m$-tuples from $B_{1}$ and $B_{2}$ and hence cannot yield the same pair of $(n+1)\times m$ sub-matrices in ${\cal B}_{1}$ and ${\cal B}_{2}$, a contradiction. Thus, $\mathbb{C}^{\prime}$ is a $(2^{k},2^{m+t};n+1,m)$-DBAC of size $\Delta$. ∎
Construction 8
Let $\mathbb{C}$ be a $(2^{k},2^{t};n,m)$-DBAC of size $\Delta$, where $3\leq m<2^{t}$, and $2\leq k,n$, whose columns are sequences with even weight. Let $A$ be any codeword in $\mathbb{C}$ and define
Let
${\cal S}=[s_{0},s_{1},\ldots,s_{2^{m-1}-1}]$
be a span $m$ half de Bruijn sequence whose first $m$ bits are zeros. Define the sequence
$B\triangleq(\overbrace{0,0,\ldots,0}^{2^{t}~{}\text{zeroes}},\overbrace{s_{1},s_{2},\ldots,s_{2^{m-1}-1},\ldots,s_{1},s_{2},\ldots,s_{2^{m-1}-1}}^{2^{t}~{}\text{copies~{}of}~{}s_{1},s_{2},\ldots,s_{2^{m-1}-1}})$
of length $2^{t}+(2^{m-1}-1)2^{t}=2^{m+t-1}$ and suppose $B=(b_{1},b_{2},\ldots,b_{2^{m+t-1}})$. Define the following codeword ${\cal B}$ in a new code $\mathbb{C}^{\prime}$: let the $i$-th column of ${\cal B}$ be ${\cal B}_{i}={\bf D}^{-1}_{b_{i}}{\cal A}_{i}$, where ${\cal A}_{i}$ is the $i$-th column of ${\cal A}$ and $1\leq i\leq 2^{m+t-1}$. Add the complement of ${\cal B}$, $\bar{{\cal B}}$, to $\mathbb{C}^{\prime}$.
The analysis for Construction 8 is very similar to that for Construction 7 as well as the proof of the following theorem whose proof is the same as that for Theorem 9.
Theorem 10
The code $\mathbb{C}^{\prime}$ defined in Construction 8 exists a $(2^{k},2^{m+t-1};n+1,m)$-DBAC of size $2\Delta$.
Construction 9
Let $\mathbb{C}$ be a $(2^{k},2^{t};n,m)$-DBAC of size $\Delta$, where $3\leq m<2^{t}$ and $2\leq n<2^{k}$, whose columns are sequences with odd weight. Let $A=[{\cal A}_{1}{\cal A}_{2}~{}\cdots~{}{\cal A}_{2^{t}}]$ be any codeword (a $2^{k}\times 2^{t}$ doubly-periodic array) of $\mathbb{C}$. Let ${\cal F}$ be a PF$(m,t)$ with no self-dual sequences, where ${\cal C}\in{\cal F}$ if and only if $\bar{{\cal C}}\notin{\cal F}$. For each pair of sequences $\{{\cal C},\bar{{\cal C}}\}$, let ${\cal S}={\cal C}$ or ${\cal S}=\bar{{\cal C}}$, where ${\cal C}\in{\cal F}$, and let
${\cal S}=[s_{1},s_{2},\ldots,s_{2^{t}}].$
For the code $\mathbb{C}^{\prime}$ define a new doubly-periodic array $B$. Let the $i$-th column of $B$ be $B_{i}={\bf D}^{-1}_{s_{i}}{\cal A}_{i}$, where $1\leq i\leq 2^{t}$. Apply the same approach on each of the $2^{t}$ distinct cyclic shifts of ${\cal S}$ to obtain $2^{t}$ codewords in $\mathbb{C}^{\prime}$.
Theorem 11
The code $\mathbb{C}^{\prime}$ defined in Construction 9 is a $(2^{k+1},2^{t};n+1,m)$-DBAC of size $2^{m-1}\Delta$.
Proof:
It is readily verified that a codeword $B$ in $\mathbb{C}^{\prime}$ is a $2^{k+1}\times 2^{t}$ array.
The perfect factor ${\cal F}$ has $2^{m-t}$ cycles and hence $2^{m-t-1}$ pairs of sequences $\{{\cal C},\bar{{\cal C}}\}$. This implies that each array $A$ is used to construct $2^{m-t-1}2^{t}=2^{m-1}$ codewords in $\mathbb{C}^{\prime}$. Since $\mathbb{C}$ is of size $\Delta$ and each array $A$ is used to construct $2^{m-t-1}2^{t}=2^{m-1}$ codewords in $\mathbb{C}^{\prime}$, it follows that $\mathbb{C}^{\prime}$ has $2^{m-1}\Delta$ codewords. Each codeword of $\mathbb{C}^{\prime}$ has $2^{k+t+1}$ distinct windows of size $(n+1)\times m$ giving total of $2^{m-1}\Delta 2^{k+t+1}$ windows of size $(n+1)\times m$ in all the arrays. Since $2^{m-1}\Delta 2^{k+t+1}=2^{nm}2^{m}=2^{(n+1)m}$ (because $\Delta 2^{k+t}=2^{nm}$), it follows that to complete the proof it is sufficient to show that each $(n+1)\times m$ array is contained at most once as an $(n+1)\times m$ window in one codeword of $\mathbb{C}^{\prime}$.
Suppose to the contrary that there exist two distinct $(n+1)\times m$ windows $X$ and $Y$ in codewords of $\mathbb{C}^{\prime}$, $B_{1}$ and $B_{2}$ (not necessarily distinct) obtained from codewords $A_{1}$ and $A_{2}$ of $\mathbb{C}$ (not necessarily distinct), respectively, for which $X=Y$. Let ${\cal X}$ and ${\cal Y}$ be the two $n\times m$ arrays obtained from $X$ and $Y$, respectively, by applying ${\bf D}$ to the columns of $X$ and $Y$, respectively (note that these columns are acyclic sequences). Clearly, if $X=Y$, then ${\cal X}={\cal Y}$. Since $A_{1}$ and $A_{2}$ are codewords in an $(2^{k},2^{t};n,m)$-DBAC, it follows that all the $n\times m$ windows in all its codewords are distinct and hence ${\cal X}={\cal Y}$ implies that they are obtained from the same position of a codeword $A$, i.e., $A_{1}=A_{2}$, using shifts of two sequences ${\cal S}_{1}$ and ${\cal S}_{2}$. Moreover, ${\cal X}={\cal Y}$ implies that they are in columns of $B_{1}$ and $B_{2}$ projected from two sets of $m$ columns associated with the same $m$-tuple ${\cal S}$ (either ${\cal S}_{1}$ or ${\cal S}_{2}$). But, this implies that $B_{1}=B_{2}$ and ${\cal S}$ is taken with the same shift to obtain $A$ and hence $X$ and $Y$ are the same window with the same position in $A$, a contradiction. Thus, $\mathbb{C}^{\prime}$ is an $(2^{k+1},2^{t};n+1,m)$-DBAC of size $2^{m-1}\Delta$. ∎
Construction 10
Let $\mathbb{C}$ be a $(2^{k},2^{t};n,m)$-DBAC of size $\Delta$, where $3\leq m<2^{t}$ and $2\leq n<2^{k}$, whose columns are sequences with odd weight. Let $A$ be any codeword in $\mathbb{C}$ and define
Let
${\cal S}=[s_{0},s_{1},\ldots,s_{2^{m-1}-1}]$
be a span $m$ half de Bruijn sequence whose first $m$ bits are zeros (therefore $[s_{1},s_{2},\ldots,s_{2^{m-1}-1}]$ is the shortened half de Bruijn sequence starting with $m-1$ zeros). Define the sequence
of length $2^{t}+(2^{m-1}-1)2^{t}=2^{m+t-1}$ and suppose $B=(b_{1},b_{2},\ldots,b_{2^{m+t-1}})$.
Define an array ${\cal B}$ in a new code $\mathbb{C}^{\prime}$. Let the $i$-th column of ${\cal B}$ be ${\cal B}_{i}={\bf D}^{-1}_{b_{i}}{\cal A}_{i}$, where ${\cal A}_{i}$ is the $i$-th column of ${\cal A}$ and $1\leq i\leq 2^{m+t-1}$.
Theorem 12
The code $\mathbb{C}^{\prime}$ defined in Construction 10 is a $(2^{k+1},2^{m+t-1};n+1,m)$-DBAC of size $\Delta$.
Proof:
A codeword ${\cal B}$ in $\mathbb{C}^{\prime}$ is a $2^{k+1}\times 2^{m+t-1}$ doubly-periodic array and hence it has $2^{k+m+t}$ windows of size $(n+1)\times m$. Since $\Delta 2^{k+m+t}=2^{m+nm}=2^{(n+1)m}$ (because $\Delta 2^{k+t}=2^{nm}$), it follows that to complete the proof it is sufficient to show that the $(n+1)\times m$ windows in the codewords of $\mathbb{C}^{\prime}$ are distinct.
The rest is the proof is very similar to that for Theorem 9 and hence is omitted. Thus, $\mathbb{C}^{\prime}$ is an $(2^{k+1},2^{m+t-1};n+1,m)$-DBAC of size $\Delta$. ∎
Each recursive construction can be applied with the codes obtained in the direct constructions or iteratively ar with codes obtained by other recursive construction. Also note that a $(2^{k},2^{t};m,n)$-DBAC is equivalent to a $(2^{t},2^{k};n,m)$-DBAC by transposing the codewords of the code. The recursive constructions can used such a code if the obtained code satisfies the requirements on the weight of the columns.
V Joining Cycles of $G_{{\cal F},m}$